OpenStudy (anonymous):

What is the limit of n to infinity of 1/(2n+2)! times by (2n)!/1

OpenStudy (anonymous):

Zero because limit of n to infinity of 1/(2n+2) = zero and zero * n =0 regardless of the other number.

OpenStudy (anonymous):

limits are not just about substituting you have to do a little algebra at first.. if there's no other way of getting a number then you substitute to infinity ... !! I don't understand well enough is this the problem we're dealing at? ... \[\LARGE \lim_{x\to \infty}\frac{1}{(2n+2)!}\cdot \frac{2n!}{1}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

0

OpenStudy (anonymous):

...I'm afraid I can't do it, better wait someone else to help you, or you can take the answer of brainshot ... but I don't know if his answer is correct or not, he didn't show up what he worked ! sorry...

OpenStudy (anonymous):

its ok thank you for trying

OpenStudy (anonymous):

..Good luck ! ;)

OpenStudy (anonymous):

You know how when 1/nx goes to infinity the number is zero, correct? That is because 1 would be so trivial when the denominator is infinity that is acts as 0. Now, we know that zero multiplied by any other number is also zero, so we know that this equation = 0.

OpenStudy (anonymous):

Ok. thanks

OpenStudy (anonymous):

np